### v1

@@ 3 Theoretical results
In this section, we analytically derive the behavior of conditional selective inference under the “marginally significant” and “highly significant” cases. For the former, we consider the behavior of $\theta(p)$ for large values of the observed $t$-statistic $x^{\text{obs}}$. For the latter, we consider the behavior of $\theta(p)$ when $x^{\text{obs}}$ is close to the critical value $\bar{x}$.
We first consider the case of marginal significance. We show that the bounds of the conditional selective confidence intervals and the median-unbiased estimator can take arbitrarily large negative values when $x^{\text{obs}}$ is close to $\bar{x}$. The following theorem is a direct application of Lemma A.3 in Kivaranovic & Leeb (2021).
@@
For $p\in(0,1)$,
$\theta(p)\to-\infty,$
as $x^{\text{obs}}\downarrow\bar{x}$,
@@
This result holds regardless of the true distribution of $X$. Its proof is entirely based on the functional form of $F(a,\theta,\bar{x})$ and does not rely on the probabilistic characteristics of the sample.
Theorem 3.1 implies that the upper bound of the conditional confidence approaches negative infinity as the estimated effect size nears the point of insignificance. Specifically, at the margin of significance, the confidence interval contains only extremely large negative values. In many practical applications, such extreme values are a-priori implausible.
Turning to the case of a highly significant parameter estimate, note that the conventional equal-tailed confidence interval at the $(1-\alpha)$-level is given by the interval $[\theta^{*}(1-\alpha/2),\theta^{*}(\alpha/2)]$ with $\theta^{*}(p)=x^{\text{obs}}-\Phi^{-1}(p)$. For example, the conventional 95% confidence interval is $[x^{\text{obs}}-1.96,x^{\text{obs}}+1.96]$. Similarly to $\theta(p)$, $\theta^{*}(p)$ can also be found by inverting a distribution function. In particular, $\theta^{*}(p)$ is the value of $\theta$ that solves the equation $p=\Phi(x^{\text{obs}}-\theta)$. The conventional point estimate is given by $\hat{\theta}=\theta^{*}(0.5)$.
@@
For $p\in(0,1)$,
$\theta(p)-\theta^{*}(p)\to 0$
as $x^{\text{obs}}\uparrow\infty$.
@@
Because $\bar{x}$ is fixed, for every $\epsilon>0$, there exists $M(\epsilon)$ such that for any $x^{\text{obs}}>M(\epsilon)$,
$\Phi(\bar{x}-(x^{\text{obs}}-\Phi^{-1}(p)))<\epsilon.$ (2)
@@
$F(x^{\text{obs}};\theta^{*}(p-\delta),\bar{x})\leq$ $\frac{\Phi(x^{\text{obs}}-\theta^{*}(p-\delta))}{1-\epsilon}$
@@
Since $F(x^{\text{obs}};\theta(p),\bar{x})=p$ and $F(x^{\text{obs}};\theta,\bar{x})$ is strictly decreasing in $\theta$ by Lemma A.2 in Kivaranovic & Leeb (2021), this establishes $\theta(p)\leq\theta^{*}(p-\delta)$ for $x^{\text{obs}}>M(\epsilon)$. Also, by inequality (2), for $x^{\text{obs}}>M(\epsilon)$,
$F(x^{\text{obs}};\theta^{*}(p+\delta),\bar{x})\geq$ $\Phi(x^{\text{obs}}-\theta^{*}(p+\delta))-\epsilon$
@@
4 Two-sided significance test
In the case of a two-sided test, the conditional selective confidence interval is not located in an extreme range. We consider the setting in Example 1 of Fithian et al. (2017) where the estimate $\hat{\theta}$ is reported only when $|x^{\text{obs}}|\geq\bar{x}$ for a critical value $\bar{x}$.
Conditional inference can be conducted similarly to the one-sided case with a different conditional distribution function. Conditionally on being reported, the distribution function of $\hat{\theta}=X$ is given by
$F_{\text{2-sided}}(a,\theta,\bar{x})=\Pr(X\leq a\mid|X|\geq\bar{x})$
$=$ $\frac{\min(\Phi(a-\theta),\Phi(-\bar{x}-\theta))+\mathbf{1}_{a>\bar{x}}(\Phi(a-\theta)-\Phi(\bar{x}-\theta))}{\Phi(-\bar{x}-\theta)+\Phi(-\bar{x}+\theta)},$
@@ 4 Two-sided significance test
where $\mathbf{1}_{a>\bar{x}}=1$ if $a>\bar{x}$ and $=0$ otherwise. Let $\theta_{\text{2-sided}}(p)$ denote the value of $\theta$ that solves $p=F_{\text{2-sided}}(x^{\text{obs}},\theta,\bar{x})$.
The median-unbiased estimator and conditional confidence interval are defined analogously to the one-sided significance case. In particular, the median-unbiased estimator is given by $\hat{\theta}_{MU,2}=\theta_{\text{2-sided}}(0.5)$ and a $(1-\alpha)$-level confidence interval is given by $[\theta_{\text{2-sided}}(1-\alpha/2),\theta_{\text{2-sided}}(\alpha/2)]$.
@@
In contrast to the one-sided significance case, $\theta_{\text{2-sided}}(p)$ converges to a finite limit as $x^{\text{obs}}\downarrow\bar{x}$. For example, at $p=0.5$, the limit solves the equation $0.5=\Phi(-\bar{x}-\theta)/(\Phi(-\bar{x}-\theta)+\Phi(-\bar{x}+\theta))$ and the median-unbiased estimate is given by $\hat{\theta}_{MU,2}=0$ as $x^{\text{obs}}\downarrow\bar{x}$.
@@
NOTE: Conditional inference after two-sided testing with critical value $\bar{x}=1.64$. The figure plots the median-unbiased estimator (solid curve) and the bounds $\theta(0.975)$ and $\theta(0.025)$ of the 95% conditional confidence interval (bold dashed curves). The thin dashed line is the 45-degree line, representing the unconditional estimate. The thin dash-dotted lines give the usual confidence interval.
Figure 2 displays the conditionally median-unbiased estimator and the upper and lower bounds of a 95% conditional selective confidence interval. The range of the horizontal axis are positive values of $x^{\text{obs}}$ that lead to a significant parameter estimate. The corresponding graphs for $x^{\text{obs}}<-\bar{x}$ are obtained by mirroring the figure around the vertical axis. The figure demonstrates several key features of conditional inference after selection based on two-sided significance.
First, the median-unbiased estimator is close to zero when the estimate is marginally significant. This stands in contrast to the one-sided significance case, where the median-unbiased estimator diverges to negative infinity.
Secondly, in the case of a highly significant parameter estimate, the conditional selective confidence interval is similar to the conventional confidence interval. This is the same result as in the one-sided significance case.
Thirdly, the lower bound of the conditional confidence interval is below the lower bound of the conventional confidence interval but not to an extreme extent. The largest difference in the lower bounds is about $0.847$, occurring at $x^{\text{obs}}\approx 2.8$.
Fourthly, the upper bound of the 95% conditional selective confidence interval falls below the conventional upper bound when $x^{\text{obs}}$ is small (for example, when it is smaller than 2.0). Very close to the critical value $\bar{x}$, the upper bound falls even below the estimate $\hat{\theta}=x^{\text{obs}}$. At this point, the conventional estimate is excluded from the conditional selective confidence interval.
Lastly, the length of the conditional selective confidence interval can be narrower than the conventional interval. Given that the upper bound is at most slightly above the conventional upper bound, the length of the conditional confidence interval is bounded and does not vastly exceed the length of the conventional confidence interval. As $x^{\text{obs}}$ approaches $\bar{x}$ from above, the length of the confidence interval decreases. The largest difference between the conditional and conventional intervals is about $2.15$ and occurs around the threshold $x^{\text{obs}}\approx\bar{x}=1.64$ where the conditional interval is the narrowest. Here, the length of the unconditional confidence interval is $3.92(=1.96\times 2)$, and that of the conditional one is $1.77$; the difference is substantial.
@@ 5 Conclusion
In this paper, we explore the properties of selective inference methods. Our main finding is that the results of these inference methods can be extreme when the parameter is only marginally significant. Therefore, applied researchers should use selective inference methods with caution. These methods should only be employed when the test statistic sufficiently exceeds the critical value.
While we consider a simple file drawer problem in this paper, results analogous to that presented in Section 3 can be obtained for general polyhedral methods, provided that the conditioning set is bounded either from below or from above. This conclusion is drawn from Lemma A.3 of Kivaranovic & Leeb (2021), which applies to general polyhedral methods. For example, conditional inference after variable selection by Lasso is affected by this issue.
In contrast, when the conditioning set is bounded, as in the case of inferences after two-sided significance testing, the problem is less pronounced. While we do not study other conditional selective inference methods such as data carving (Fithian et al., 2017) and selection with a randomized response (Tian & Taylor, 2018) in this paper, these methods yield confidence intervals whose expected lengths are bounded (Kivaranovic & Leeb, 2024). To the best of our knowledge, the location characteristics of their confidence intervals have not yet been investigated in the literature. We conjecture that the problem manifests less dramatically with these methods.

### v2

@@ 2 Theoretical results
In this section, we analytically derive the behavior of conditional selective inference under the “marginally significant” and “highly significant” cases. For the former, we consider the behavior of $\theta(p)$ when $x^{\text{obs}}$ is close to the critical value $c$. For the latter, we consider the behavior of $\theta(p)$ for large values of the observed $t$-statistic $x^{\text{obs}}$.
For the case of marginal significance, Lemma A.3 in Kivaranovic & Leeb (2021) implies that, as $x^{\text{obs}}$ approaches $c$, the conditional confidence interval and the median-unbiased estimator drift toward negative infinity.
@@
For $p\in(0,1)$, as $x^{\text{obs}}\downarrow c$
$\theta(p)\to-\infty.$
@@
This result holds regardless of the true distribution of $X$. Its proof is entirely based on the functional form of $F(a,\theta,c)$ and does not rely on the probabilistic characteristics of the sample.
Theorem 2.1 implies that, as the estimated effect size approaches insignificance, every possible parameter value is ruled out eventually. This is formalized in the following corollary:
Corollary 2.1.
For any $\theta\in\mathbb{R}$ and $\alpha\in(0,0.5)$, there exists $\delta>0$ such that if $c<x^{\text{obs}}<c+\delta$, then $\theta\notin[\theta(1-\alpha/2),\theta(\alpha/2)]$.
For a scenario where one has information or beliefs about a plausible range of parameter values, this means that, if $x^{\text{obs}}$ realizes too close to the margin of significance, the conditional confidence interval will not contain any plausible values.
Turning to the case of a highly significant parameter estimate, note that the conventional equal-tailed confidence interval at the $(1-\alpha)$-level is given by the interval $\theta^{*}(1-\alpha/2),\theta^{*}(\alpha/2)]$ with $\theta^{*}(p)=x^{\text{obs}}-\Phi^{-1}(p)$. For example, the conventional 95% confidence interval is $x^{\text{obs}}-1.96,x^{\text{obs}}+1.96]$. Similarly to $\theta(p)$, $\theta^{*}(p)$ can also be found by inverting a distribution function. In particular, $\theta^{*}(p)$ is the value of $\theta$ that solves the equation $p=\Phi(x^{\text{obs}}-\theta)$. The conventional point estimate is given by $\hat{\theta}=\theta^{*}(0.5)$.
@@
For $p\in(0,1)$, as $x^{\text{obs}}\uparrow\infty$,
$\theta(p)-\theta^{*}(p)\to 0.$
@@
Because $c$ is fixed, for every $\epsilon>0$, there exists $M(\epsilon)$ such that for any $x^{\text{obs}}>M(\epsilon)$,
$\Phi(c-(x^{\text{obs}}-\Phi^{-1}(p)))<\epsilon.$ (2)
@@
$F(x^{\text{obs}};\theta^{*}(p-\delta),c)\leq$ $\frac{\Phi(x^{\text{obs}}-\theta^{*}(p-\delta))}{1-\epsilon}$
@@
Since $F(x^{\text{obs}};\theta(p),c)=p$ and $F(x^{\text{obs}};\theta,c)$ is strictly decreasing in $\theta$ by Lemma A.2 in Kivaranovic & Leeb (2021), this establishes $\theta(p)\leq\theta^{*}(p-\delta)$ for $x^{\text{obs}}>M(\epsilon)$. Also, by inequality (2), for $x^{\text{obs}}>M(\epsilon)$,
$F(x^{\text{obs}};\theta^{*}(p+\delta),c)\geq$ $\Phi(x^{\text{obs}}-\theta^{*}(p+\delta))-\epsilon$
@@
3 Alternative scenarios
We now discuss two variations of the file drawer problem: file drawer with randomized response and file drawer with two-sided testing. In these alternative settings, it is possible to construct conditional inference procedures that avoid the location problem.
3.1 The randomized file drawer problem
We first consider a modification of the file drawer problem that redefines significance to depend on an additional randomization. Selective inference in this setting falls within the framework of “selection with a randomized response” that has garnered attention in the recent statistics literature (Fithian et al., 2017; Tian & Taylor, 2018; Panigrahi et al., 2021; Panigrahi & Taylor, 2023). The randomization is expected to “smooth out” erratic behavior close to the threshold of significance.
To describe the randomized file drawer problem, let $X\sim N(\theta,1)$ as before and introduce a randomization $W$ that is independent of $X$ and satisfies $W\sim N(0,\eta^{2})$. The parameter estimate $\hat{\theta}=X$ is “significant after randomization” if $X+W\geq c$ for a critical value $c$. The parameter $\eta$ is a tuning parameter that controls the degree of randomization.
An example of a randomization mechanism is data carving (Fithian et al., 2017), that is, using a random subset of the data for selection and the entire sample for post-selection inference. In particular, suppose that we have a random sample $\{Z_{i}\}_{i=1}^{n}$, where $Z_{i}\sim N(\theta,n)$. The first $n_{1}$ observations determine the selection. Let
$X=\frac{1}{n}\sum_{i=1}^{n}Z_{i},\quad W=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}Z_{i}-\frac{1}{n}\sum_{i=1}^{n}Z_{i}.$
In this case, $X+W=\sum_{i=1}^{n_{1}}Z_{i}/n_{1}$. Note that $X\sim N(\theta,1)$ and $X+W\sim N(\theta,n/n_{1})$ so that $\eta^{2}=n/n_{1}-1$. To clarify the role of $\eta$, note that setting $n_{1}=n/2$ (using half the sample for selection) gives $\eta^{2}=1$, while setting $n_{1}=n$ (using the entire sample for both selection and inference) gives $\eta^{2}=0$. Selection and inference are closely related when $\eta$ is small, and they become more independent as $\eta$ increases.
To conduct statistical inference on the parameter of interest $\theta$, consider the conditional probability
$F^{R}(a;\theta,c)=P(X\leq a\mid X+W\geq c).$
The only difference to (1) is that the conditioning set now accounts for the randomization. Solve
$p=F^{R}(x^{\text{obs}};\theta^{R}(p),c),$
for $\theta^{R}(p)$. The conditionally median-unbiased estimator is $\theta^{R}(0.5)$, and a $100(1-\alpha)\%$ conditional confidence interval is $\theta^{R}(1-\alpha/2),\theta^{R}(\alpha/2)]$. [^1] Panigrahi & Taylor (2023) consider the maximum likelihood estimation based on the conditional likelihood. In particular, Section 2 of their paper considers the file drawer problem.
Figure 2: Conditional inferences based on selection with a randomized response
NOTE: Conditional inference based on one-sided testing with a randomized response with $\eta^{2}=1$ and $c=1.64$. The figure plots the median-unbiased estimator (solid curve) and the bounds $\theta(0.975)$ and $\theta(0.025)$ of the 95% conditional confidence interval (dashed curves). The thin dashed line is the 45-degree line, representing the unconditional estimate. The shaded region represents the conventional unconditional 95% confidence interval.
Figure 2 shows that selective inference with a randomized response avoids the location problem observed in the original file-drawer problem: both the median-unbiased estimator and the conditional confidence interval remain “stable,” in the sense that they do not diverge rapidly toward negative infinity, as $x^{\text{obs}}$ approaches the margin of significance.
The use of the randomized-response approach entails restrictions on the significance test. For example, in the case of data-carving, it requires that a substantial proportion of the data is set aside when testing significance. This corresponds to choosing a sufficiently large tuning parameter $\eta$. If $\eta$ is small, the inference results become similar to those of the original polyhedral method, potentially reintroducing the location problem. Testing significance with only part of the sample is suboptimal for empirical studies that focus on establishing significance and consider effect size only as a secondary concern. It is also not suitable for reexamining effect size in existing empirical studies that have already obtained significance results based on the entire sample.
Randomized response methods shine in empirical designs where effect size is the primary focus and where the selection procedure can be designed appropriately from the outset. For example, in high-dimensional settings where selection is employed primarily for dimension reduction — using methods such as the Lasso — the selection stage is not itself the central focus. In these cases, randomized-response techniques, including randomized Lasso (Panigrahi & Taylor, 2023), can be especially advantageous.
3.2 The two-sided file drawer problem
In our original file drawer problem, the estimate is reported when it is “significantly positive”. We now consider the alternative where estimates are reported if they are “significantly different from zero” (see Example 1 in Fithian et al. (2017)). Formally, we select estimates that satisfy $|x^{\text{obs}}|\geq c$ for a critical value $c$.
Polyhedral inference can be implemented analogously to the one-sided case based on the conditional distribution function
$F_{\text{2-sided}}(a,\theta,c)=\Pr(X\leq a\mid|X|\geq c)$
$=$ $\frac{\min(\Phi(a-\theta),\Phi(-c-\theta))+\mathbf{1}_{a>c}(\Phi(a-\theta)-\Phi(c-\theta))}{\Phi(-c-\theta)+\Phi(-c+\theta)},$
@@ 3.2 The two-sided file drawer problem
where $\mathbf{1}_{a>c}=1$ if $a>c$ and =0=0 otherwise. Let $\theta_{\text{2-sided}}(p)$ denote the value of $\theta$ that solves $p=F_{\text{2-sided}}(x^{\text{obs}},\theta,c)$. The median-unbiased estimator is given by $\hat{\theta}_{MU,2}=\theta_{\text{2-sided}}(0.5)$ and a $(1-\alpha)$-level confidence interval is given by $\theta_{\text{2-sided}}(1-\alpha/2),\theta_{\text{2-sided}}(\alpha/2)]$.
@@
In contrast to the one-sided case, $\theta_{\text{2-sided}}(p)$ converges to a finite limit as $x^{\text{obs}}\downarrow c$. For example, at $p=0.5$, the limit solves the equation $0.5=\Phi(-c-\theta)/(\Phi(-c-\theta)+\Phi(-c+\theta))$ and the median-unbiased estimate is given by $\hat{\theta}_{MU,2}=0$ as $x^{\text{obs}}\downarrow c$.
@@
NOTE: Conditional inference after two-sided testing with critical value $c=1.64$ with and without randomized response ($\eta^{2}=1$). The dashed and dotted lines give the confidence intervals with and without randomized response, respectively. The solid and dash-dotted lines give the corresponding median-unbiased estimates. The thin dashed line is the 45-degree line, representing the unconditional estimate. The shaded region represents the conventional unconditional 95% confidence interval.
Figure 3 displays the polyhedral conditional confidence interval and the corresponding median-unbiased estimate. For comparison, it also shows the confidence interval and point estimate based on a two-sided variant of the model with a randomized response discussed in Section 3.1. The horizontal range of the figure are positive values of the $t$-statistic $x^{\text{obs}}$ for which we reject. A corresponding figure for negative values of $x^{\text{obs}}$ is obtained by reflecting the figure around the vertical axis.
Even without introducing randomization, the two-sided approach avoids the location problem observed in the one-sided case. For marginal significance, the conditional median-unbiased estimator and the conditional confidence interval no longer exhibit the extreme values that arise in the one-sided setting. In particular, the median-unbiased estimator converges to zero as we approach insignificance and the lower bound of the conditional confidence interval lies slightly below the lower bound of the conventional confidence interval. At most, it lies $0.847$ standard deviations below the conventional lower bound, which occurs at $x^{\text{obs}}\approx 2.8$. By contrast, in the one-sided case without randomization, both the median-unbiased estimator and the lower bound of the confidence interval diverge to negative infinity as the $t$-statistic approaches the critical value from above.
Regardless of whether the selection contains a randomized response, the upper bound of the conditional confidence interval after two-sided selection lies below the conventional upper bound when $x^{\text{obs}}$ is small (for example, for values less than $2.0$). Approaching the critical value from above, the gap to the conventional upper bound becomes especially pronounced in the non-randomized case; eventually, it drops below the conventional estimator $\hat{\theta}=x^{\text{obs}}$, producing a confidence interval around zero whose length approaches only $1.77$ standard deviations. The conventional interval has length $2\times 1.96=3.92$ standard deviations. Hence, conditional polyhedral inference can be substantially more precise than conventional inference. This stands in sharp contrast to the one-sided case, where polyhedral inference produces very wide confidence intervals at the margin of significance.
Comparing conditional intervals after selection with or without randomization, neither approach yields a uniformly tighter confidence interval. As discussed above, without randomization the conditional confidence interval contracts sharply at the margin of significance. By contrast, with randomization the interval is wider and more closely aligned with the unconditional interval. At moderate values of the $t$-statistic, however, the randomized interval is shorter than its non-randomized counterpart. The difference is largest at $0.44$ standard deviations, which occurs when $x^{\text{obs}}\approx 2.76$. From a practical standpoint, therefore, neither interval provides a decisive advantage, even in settings where both approaches are empirically viable.
@@ 4 Conclusion
In this paper, we examine the properties of selective inference methods. Our principal finding is that these methods can infer values or value ranges that are highly negative when the parameter of interest is only marginally significant. In many practical applications, such extreme values are ruled out by a-priori information about the parameter space. In this case, selective inference cannot infer any plausible parameter values.
While we consider a simple file drawer problem in this paper, results analogous to that presented in Section 2 can be obtained for general polyhedral methods, provided that the conditioning set is bounded either from below or from above. This is implied by Lemma A.3 of Kivaranovic & Leeb (2021), which applies to general polyhedral methods. For example, conditional inference after variable selection by the Lasso is affected by this issue.
Our main result covers the case of selection by a one-sided test. We demonstrate that selection with a randomized response and selection based on a two-sided test guard against the location problem at the cost of redefining the selection mechanism.